# Porosity and supercyclic operators on solid Banach function spaces

### v6

@@ 2. Preliminaries
Indeed, since $T$ is topologically semi-transitive, we can find some $n_{1}\in\mathbb{N}$ and some $\lambda_{1}\in\mathbb{C}\setminus\{0\}$ such that $\lambda_{1}T^{n_{1}}(O_{1})\cap O_{2}\neq\varnothing$ Now, since $\lambda_{1}T^{n_{1}}$ is invertible, it is an open map, hence $\lambda_{1}T^{n_{1}}(O_{1})$ is open. Therefore, there exists some $\tilde{n}_{2}\in\mathbb{N}$ and some $\tilde{\lambda}_{2}\in\mathbb{C}\setminus\{0\}$ such that $\tilde{\lambda}_{2}T^{{\tilde{n}}_{2}}(\lambda_{1}T^{n_{1}}(O_{1}))\cap O_{2}\neq\varnothing.$ Put $n_{2}=n_{1}+\tilde{n}_{2}$ and $\lambda_{2}=\lambda_{1}\tilde{\lambda}_{2}.$ Proceeding inductively, we can construct the desired sequences $\{n_{k}\}_{k}$ and $\{\lambda_{k}\}_{k}.$
@@ 3. Supercyclicity of weighted composition operators
$\lim_{x\rightarrow\infty}\left\|\chi_{K\backslash E_{k}}\right\|_{\mathcal{F}}=0$
@@
$(\prod_{j=0}^{n_{k}-1}(w\circ\alpha^{j})(x)))^{-1}<\frac{|\lambda_{k}|\text{ }}{{2^{k}}-1}$
@@
Next we prove $ii)\Rightarrow i).$ Let $\mathcal{O}_{1},$ and $\mathcal{O}_{2}$ be non-empty open subsets of $\mathcal{F}.$ Then we can find some $f\in(\mathcal{O}_{1}\setminus\{0\})\cap\mathcal{F}_{b_{c}}$ and $g\in(\mathcal{O}_{2}\setminus\{0\})\cap\mathcal{F}_{b_{c}}$ since $\mathcal{O}_{1}\setminus\{0\}),\mathcal{O}_{2}\setminus\{0\})$ are also open, non-empty and $\mathcal{F}_{b_{c}}$ is dense in $\mathcal{F}.$ As in the proof of [6, Theorem 1], set $K=supp\text{ }f\cup supp\text{ }g.$ Choose the strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ and the sequence of Borel subsets $\{E_{k}\}_{k}$ of $K$ that satisfy the assumptions of $ii)$ with respect to $K.$ As shown in the proof of [6, Theorem 1], we have that $\parallel f-f\chi_{E_{k}}\parallel_{\mathcal{F}}\rightarrow 0$ and $\parallel g-g\chi_{E_{k}}\parallel_{\mathcal{F}}\rightarrow 0$ when $k\rightarrow\infty,$ so we may without loss of generality, asume that $f\chi_{E_{k}}\in\mathcal{O}_{1}\setminus\{0\}$ and $g\chi_{E_{k}}\in\mathcal{O}_{2}\setminus\{0\}$ for all $k.$ Therefore, $\tilde{T}_{\alpha,w}^{n_{k}}(f\chi_{E_{k}})\neq 0$ and $\tilde{S}_{\alpha,w}^{n_{k}}(g\chi_{E_{k}})\neq 0$ for all $k$ because $\tilde{T}_{\alpha,w}$ and $\tilde{S}_{\alpha,w}$ are invertible. As in the proof of [6, Theorem 1], we have for each $k\in\mathbb{N}$ that
@@
By combining (3) and (4) together with the assumptions in $ii),$ it is not hard to deduce that $v_{k}\rightarrow f$ and
@@
$\lim_{x\rightarrow\infty}\left\|\chi_{K\backslash E_{k}}\right\|_{\mathcal{F}}=0$
@@
Let Ω=, $\Omega=\mathbb{R},\text{ }\alpha:\mathbb{R}\rightarrow\mathbb{R}$ be given by $\alpha(t)=t-1$ for all $t\in\mathbb{R}$ and $w$ be a continuous bounded positive function on $\mathbb{R}.$ If there exist some $M,\delta,K_{1},K_{2}>0$ such that $1<M-\delta\leq w(t)\leq M$ for all $t\leq-K_{1}$ and $w(t)=1$ for all $t\geq K_{2},$ then the conditions of Proposition 3.6 and Proposition 3.11 are satisfied.
@@
[…]
@@
[…]
@@ 4. Supercyclic adjoints of weighted composition operators
[…]
@@
Next ve prove $ii)\Rightarrow i).$ Given two non-empty open subsets $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ of $M(\Omega),$ since $\mathcal{O}_{1}\setminus\{0\}$ and $\mathcal{O}_{2}\setminus\{0\}$ are also open and non-empty, we can find some $\mu\in\mathcal{O}_{1}\setminus\{0\}$ and some $v\in\mathcal{O}_{2}\setminus\{0\}$ such that $|\mu|(K^{c})=|v|(K^{c})=0$ for some compact subset $K$ of $\Omega.$ Choose the strictly increasing sequence $\{n_{k}\}_{k}\in\mathbb{N},$ and the sequences $\{A_{k}\}_{k},\{B_{k}\}_{k}$ of Borel subsets of $K$ satisfying the assumptions of $ii)$ with respect to $\mu,v$ and $K.$ For each $k\in\mathbb{N},$ let $\tilde{\mu_{k}},\tilde{v_{k}}$ be the measures in $M(\Omega)$ given by $\tilde{\mu_{k}}(E)=\mu(E\cap A_{k}^{c})$ and $\tilde{v_{k}}(E)=v(E\cap B_{k}^{c})$ for every measurable subset $E$ of $\Omega.$ By the same arguments as in the proof of [18, Proposition 3.1], we can deduce that for all $k\in\mathbb{N}$ it holds that
@@
$\eta_{k}=\tilde{\mu_{k}}+\frac{\parallel\tilde{T}_{\alpha,w}^{*n_{k}}(\tilde{\mu_{k}})\parallel^{\frac{1}{2}}}{\parallel\tilde{S}_{\alpha,w}^{*n_{k}}(\tilde{v_{k}})\parallel^{\frac{1}{2}}}\tilde{S}_{\alpha,w}^{*n_{k}}(\tilde{v_{k}}).$
By combining $(1)$ and $(2)$ together with the arguments from the proof of the implication $ii)\Rightarrow i)$ in Proposition 3.5, we can deduce from the assumptions in $ii)$ that $\eta_{k}\rightarrow\mu$ and
@@
Let Ω=, $\Omega=\mathbb{R},\text{ }\alpha:\mathbb{R}\rightarrow\mathbb{R}$ be given by $\alpha(t)=t+1$ for all $t\in\mathbb{R}$ and $w$ be a continuous bounded positive function on $\mathbb{R}.$ If there exist some $M,\delta,K_{1},K_{2}>0$ such that $1<M-\delta\leq w(t)\leq M$ for all $t\leq-K_{1}$ and $w(t)=1$ for all $t\geq K_{2},$ then the conditions of Proposition 4.2 are satisfied, however, by [18, Proposition 3.1] it follows that $\tilde{T}_{\alpha,w}^{*}$ is not topologically transitive on $M(\Omega).$ On the other hand, if $\alpha(t)=t-1$ for all $t\in\mathbb{R},\frac{1}{M}\leq w(t)\leq\frac{1}{M-\delta}$ for all $t\leq-K_{1}$ and $1=w(t)$ for all $t\geq K_{2},$ then the conditions of Proposition 4.1 are satisfied, however, $\tilde{T}_{\alpha,w}^{*}$ is neither topologically Cesáro hyper-transitive nor topologically transitive on $M(\Omega),$ which follows from Proposition 4.2 and [18, Proposition 3.1].
@@ 5. Porosity and weighted composition operators
Let $10<\lambda<$ 1. A subset $E$ of a metric space $X$ is called $\lambda$-porous at $x\in E$ if for each $\delta>0$ there is an element $y\in B(x;\delta)\setminus\{x\}$ such that
@@
:=$\mathcal{F}:=\big{\{}\Gamma_{g}:\,\,g\in C_{0}(\mathbb{R})\big{\}}.$
@@
$h(t)=\begin{cases}g(t)+\delta\text{ for }t\in[-N,N],\\ \beta^{-1}g(t)\text{ for }t\in(-\infty,-N-1]\cup[N+1,\infty),\\ \beta^{-1}g(t-N-1)+(t+N+1)\left(\delta+g(-N)-\beta^{-1}g(-N-1)\right),\text{ for }t\in(-N-1,-N),\\ g(N)+\delta+(t-N)\left(\beta^{-1}g(N+1)-\delta-g(N)\right)\text{ for }t\in(N,N+1).\end{cases}$
Then $h\in\Gamma_{g}.$ To simplify notation, we set $M=2N$ and for each $j\in\mathbb{Z}\cap[-N,N]$, we put $x_{j}=j-N$ for $j\in\{0,1,\dots,M\}.$ We let then $\eta$ be the function from $\{x_{0},x_{1},\dots,x_{M}\}$ into $\mathbb{R}$ given by $\eta(x_{i})=\begin{cases}\frac{k(x_{i})}{|k(x_{i})|}&\text{if }k(x_{i})\neq 0,\ i\in\{0,1,\dots,M\},\\ 1&\text{if }k(x_{i})=0,\ i\in\{0,1,\dots,M\}.\end{cases}$ Further, we let $\tilde{\eta}$ be the piecewise linear function on $[-N,N]$ connecting the points $\left(x,\eta(x_{i})\right)$ where $i\in\{0,\dots,M\}.$ More precisely, on each segment $[x_{i-1},x_{i}]$ with $i\in\{1,\dots,M\},$ the function $\tilde{\eta}$ is given by $\tilde{\eta}(t)=\eta(x_{i-1})+\frac{t-x_{i-1}}{x_{i}-x_{i-1}}\left(\eta(x_{i})-\eta(x_{i-1})\right)$ for $t\in[x_{i-1},x_{i}].$
@@
$\mathcal{E}(t)=\begin{cases}k(t)+\delta\tilde{\eta}(t)\text{ for }t\in[-N,N],\\ h(t)\text{ for }t\in(-\infty,-N-1]\cup[N+1,\infty),\\ h(-N-1)+(t+N+1)\left(k(-N)+\delta\tilde{\eta}(-N)-h(-N-1)\right)\text{ for }t\in(-N-1,-N),\\ k(N)+\delta\tilde{\eta}(N)+(t-N)\left(h(N+1)-k(N)-\delta\tilde{\eta}(N)\right)\text{ for }t\in(N,N+1).\end{cases}.$ Then $\mathcal{E}\in C_{0}(\mathbb{R})$ since $h\in C_{0}(\mathbb{R}),$ and $k,\tilde{\eta}$ are continuous. We notice that by the triangle inequality, for all $t\in(N,N+1)$ we have
@@
into $\mathbb{C}$ given by $\Theta(m)=\begin{cases}\frac{v(m)}{|v(m)|}&\text{if }v(m)\neq 0,\\ 1&\text{if }v(m)=0,\end{cases}$ and we let $\tilde{\Theta}$ be the piecewise linear function from $(-\infty,-N-1]\cup[N+1,\infty)$ into $\mathbb{C}$ connecting the points $(m,\Theta(m))$ where $m\in\mathbb{Z}\cap\left((-\infty,-N-1]\cup[N+1,\infty)\right).$ More precisely, on each segment $[m-1,m]$ with $m\in\mathbb{Z}\cap((-\infty,-N-1]\cup[N+2,\infty)),$ the function $\tilde{\Theta}$ is given as
@@
$\gamma(t)=\begin{cases}v(t)\text{ if }t\in[-N,N],\\ v(t)+\beta|u(t)-v(t)|\tilde{\Theta}(t)\text{ for }t\in(-\infty,-N-1]\cup[N+1,\infty),\\ v(t)+(t-N)|\beta|u(N+1)-v(N+1)|\tilde{\Theta}(N+1)\text{ for }t\in(N,N+1),\\ v(t)-(t+N)\beta|u(-N-1)-v(-N-1)|\tilde{\Theta}(-N-1)\text{ for }t\in(-N-1,-N).\end{cases}$ Since $\tilde{\Theta}$ is continuous, $|\tilde{\Theta}(t)|\leq 1$ for all $t\in(-\infty,-N-1]\cup[N+1,\infty)$ and $u,v\in C_{0}(\mathbb{R}),$ it is not hard to see that $\gamma\in C_{0}(\mathbb{R}).$ Also, from the construction of the function $\gamma$ it follows that $\|\gamma-v\|_{\infty}\leq\beta\|u-v\|_{\infty}\leq\lambda\|u-v\|_{\infty}.$ For $m\in\mathbb{Z}\cap\left((-\infty,-N-1]\cup[N+1,\infty)\right)$ we have
@@
Consider the weighted composition operator $\tilde{T}_{\alpha,w}$ on $C_{0}(\mathbb{R})$ given by $\tilde{T}_{\alpha,w}(f)=w\cdot(f\circ\alpha),$ where $0<w,w^{-1}\in C_{b}(\mathbb{R}),$ and $\alpha$ is a homeomorphism of $\mathbb{R}.\$ If $\lim_{n\to\infty}\prod_{k=1}^{n}(w\circ\alpha^{-k})^{-1}(n)=0,$ then the set
$\left\{f\in C_{0}(\mathbb{R}):\|\tilde{T}_{\alpha,w}^{n}(f)\|_{\infty}\geq 1\text{ for all }n\in\mathbb{N}\right\}\$
is not $\sigma\text{-porous. In particular, the set of non-hypercyclic vectors for theoperator }\tilde{T}_{\alpha,w}$ is not $\sigma\text{-porous in }C_{0}(\mathbb{R}).$
@@
$+(t-n)\left(\prod_{k=1}^{n+1}(w\circ\alpha^{-k})^{-1}(n+1)-\prod_{k=1}^{n}(w\circ\alpha^{-k})^{-1}(n)\right)$
[…]

### v7

@@ 2. Preliminaries
Indeed, since $T$ is topologically semi-transitive, we can find some $n_{1}\in\mathbb{N}$ and some $\lambda_{1}\in\mathbb{C}\setminus\{0\}$ such that $\lambda_{1}T^{n_{1}}(O_{1})\cap O_{2}\neq\varnothing$ . Now, since $\lambda_{1}T^{n_{1}}$ is invertible, it is an open map, hence $\lambda_{1}T^{n_{1}}(O_{1})$ is open. Therefore, there exists some $\tilde{n}_{2}\in\mathbb{N}$ and some $\tilde{\lambda}_{2}\in\mathbb{C}\setminus\{0\}$ such that $\tilde{\lambda}_{2}T^{{\tilde{n}}_{2}}(\lambda_{1}T^{n_{1}}(O_{1}))\cap O_{2}\neq\varnothing.$ Put $n_{2}=n_{1}+\tilde{n}_{2}$ and $\lambda_{2}=\lambda_{1}\tilde{\lambda}_{2}.$ Proceeding inductively, we can construct the desired sequences $\{n_{k}\}_{k}$ and $\{\lambda_{k}\}_{k}.$
@@ 3. Supercyclicity of weighted composition operators
$lim_{x\rightarrow\infty}\left\|\chi_{K\backslash E_{k}}\right\|_{\mathcal{F}}=0$
@@
[…]
@@
Next we prove $ii)\Rightarrow i).$ Let $\mathcal{O}_{1},$ and $\mathcal{O}_{2}$ be non-empty open subsets of $\mathcal{F}.$ Then we can find some $f\in(\mathcal{O}_{1}\setminus\{0\})\cap\mathcal{F}_{b_{c}}$ and $g\in(\mathcal{O}_{2}\setminus\{0\})\cap\mathcal{F}_{b_{c}}$ since $\mathcal{O}_{1}\setminus\{0\}),\mathcal{O}_{2}\setminus\{0\})$ are also open, non-empty and $\mathcal{F}_{b_{c}}$ is dense in $\mathcal{F}.$ As in the proof of [6, Theorem 1], set $K=supp\text{ }f\cup supp\text{ }g.$ Choose the strictly increasing sequence $\{n_{k}\}_{k}\subseteq\mathbb{N}$ and the sequence of Borel subsets $\{E_{k}\}_{k}$ of $K$ that satisfy the assumptions of ii)ii) with respect to $K.$ As shown in the proof of [6, Theorem 1], we have that $\parallel f-f\chi_{E_{k}}\parallel_{\mathcal{F}}\rightarrow 0$ and $\parallel g-g\chi_{E_{k}}\parallel_{\mathcal{F}}\rightarrow 0$ when $k\rightarrow\infty,$ so we may without loss of generality, asume that $f\chi_{E_{k}}\in\mathcal{O}_{1}\setminus\{0\}$ and $g\chi_{E_{k}}\in\mathcal{O}_{2}\setminus\{0\}$ for all $k.$ Therefore, $\tilde{T}_{\alpha,w}^{n_{k}}(f\chi_{E_{k}})\neq 0$ and $\tilde{S}_{\alpha,w}^{n_{k}}(g\chi_{E_{k}})\neq 0$ for all $k$ because $\tilde{T}_{\alpha,w}$ and $\tilde{S}_{\alpha,w}$ are invertible. As in the proof of [6, Theorem 1], we have for each $k\in\mathbb{N}$ that
@@
By combining (3) and (4) together with the assumptions in ii),ii), it is not hard to deduce that $v_{k}\rightarrow f$ and
@@
$lim_{x\rightarrow\infty}\left\|\chi_{K\backslash E_{k}}\right\|_{\mathcal{F}}=0$
@@
Let $\Omega=\mathbb{R},\text{ }\alpha:\mathbb{R}\rightarrow\mathbb{R}$ be given by $\alpha(t)=t-1$ for all $t\in\mathbb{R}$ and $w$ be a continuous bounded positive function on $\mathbb{R}.$ If there exist some $M,\delta,K_{1},K_{2}>0$ such that $1<M-\delta\leq w(t)\leq M$ for all $t\leq-K_{1}$ and $w(t)=1$ for all $t\geq K_{2},$ then the conditions of Proposition 3.6 and Proposition 3.11 are satisfied.
@@
$\int\left|W^{-n}\left(f\right)\right|^{2}d\mu\leq\sup_{t\in\left[-m,m\right]}(\prod^{n-1}_{j=0}(w\circ\alpha^{j})^{-1}(t))^{2}\left\|f\right\|_{2}^{2}$
@@
$\int\left|W^{n}\left(f\right)\right|^{2}d\mu\leq\sup_{t\in\left[-m,m\right]}(\prod^{n-1}_{j=0}(w\circ\alpha^{j-n})(t))^{2}\left\|f\right\|_{2}^{2}.$
@@ 4. Supercyclic adjoints of weighted composition operators
$\tilde{T}_{\alpha,w}^{*}(\mu)(E)=\int_{E}w\circ\alpha^{-1}\,d\mu\ \circ\alpha^{-1}$
@@
Next ve prove $ii)\Rightarrow i).$ Given two non-empty open subsets $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ of $M(\Omega),$ since $\mathcal{O}_{1}\setminus\{0\}$ and $\mathcal{O}_{2}\setminus\{0\}$ are also open and non-empty, we can find some $\mu\in\mathcal{O}_{1}\setminus\{0\}$ and some $v\in\mathcal{O}_{2}\setminus\{0\}$ such that $|\mu|(K^{c})=|v|(K^{c})=0$ for some compact subset $K$ of $\Omega.$ Choose the strictly increasing sequence $\{n_{k}\}_{k}\in\mathbb{N},$ and the sequences $\{A_{k}\}_{k},\{B_{k}\}_{k}$ of Borel subsets of $K$ satisfying the assumptions of ii)ii) with respect to $\mu,v$ and $K.$ For each $k\in\mathbb{N},$ let $\tilde{\mu_{k}},\tilde{v_{k}}$ be the measures in $M(\Omega)$ given by $\tilde{\mu_{k}}(E)=\mu(E\cap A_{k}^{c})$ and $\tilde{v_{k}}(E)=v(E\cap B_{k}^{c})$ for every measurable subset $E$ of $\Omega.$ By the same arguments as in the proof of [18, Proposition 3.1], we can deduce that for all $k\in\mathbb{N}$ it holds that
@@
[…]
By combining $(1)$ and $(2)$ together with the arguments from the proof of the implication $ii)\Rightarrow i)$ in Proposition 3.5, we can deduce from the assumptions in ii)ii) that $\eta_{k}\rightarrow\mu$ and
@@
Let $\Omega=\mathbb{R},\text{ }\alpha:\mathbb{R}\rightarrow\mathbb{R}$ be given by $\alpha(t)=t+1$ for all $t\in\mathbb{R}$ and $w$ be a continuous bounded positive function on $\mathbb{R}.$ If there exist some $M,\delta,K_{1},K_{2}>0$ such that $1<M-\delta\leq w(t)\leq M$ for all $t\leq-K_{1}$ and $w(t)=1$ for all $t\geq K_{2},$ then the conditions of Proposition 4.2 are satisfied, however, by [18, Proposition 3.1] it follows that $\tilde{T}_{\alpha,w}^{*}$ is not topologically transitive on $M(\Omega).$ On the other hand, if $\alpha(t)=t-1$ for all $t\in\mathbb{R},\frac{1}{M}\leq w(t)\leq\frac{1}{M-\delta}$ for all $t\leq-K_{1}$ and $1=w(t)$ for all $t\geq K_{2},$ then the conditions of Proposition 4.1 are satisfied, however, $\tilde{T}_{\alpha,w}^{*}$ is neither topologically Cesáro hyper-transitive nor topologically transitive on $M(\Omega),$ which follows from Proposition 4.2 and [18, Proposition 3.1].
@@ 5. Porosity and weighted composition operators
Let $0<\lambda<1$. A subset $E$ of a metric space $X$ is called $\lambda$-porous at $x\in E$ if for each $\delta>0$ there is an element $y\in B(x;\delta)\setminus\{x\}$ such that
@@
$\mathcal{F}:=\big\{\Gamma_{g}:\,\,g\in C_{0}(\mathbb{R})\big\}.$
@@
[…]
Then $h\in\Gamma_{g}.$ To simplify notation, we set $M=2N$ and for each $j\in\mathbb{Z}\cap[-N,N]$, we put $x_{j}=j-N$ for $j\in\{0,1,\dots,M\}.$ We let then $\eta$ be the function from $\{x_{0},x_{1},\dots,x_{M}\}$ into $\mathbb{R}$ given by $\eta(x_{i})=\begin{cases}\frac{k(x_{i})}{|k(x_{i})|}&\text{if }k(x_{i})\neq 0,\ i\in\{0,1,\dots,M\},\\ 1&\text{if }k(x_{i})=0,\ i\in\{0,1,\dots,M\}.\end{cases}$ Further, we let $\tilde{\eta}$ be the piecewise linear function on [N,N][-N,N] connecting the points $\left(x,\eta(x_{i})\right)$ where $i\in\{0,\dots,M\}.$ More precisely, on each segment $x_{i-1},x_{i}]$ with $i\in\{1,\dots,M\},$ the function $\tilde{\eta}$ is given by $\tilde{\eta}(t)=\eta(x_{i-1})+\frac{t-x_{i-1}}{x_{i}-x_{i-1}}\left(\eta(x_{i})-\eta(x_{i-1})\right)$ for $t\in[x_{i-1},x_{i}].$
@@
[…]
@@
into $\mathbb{C}$ given by $\Theta(m)=\begin{cases}\frac{v(m)}{|v(m)|}&\text{if }v(m)\neq 0,\\ 1&\text{if }v(m)=0,\end{cases}$ and we let $\tilde{\Theta}$ be the piecewise linear function from $(-\infty,-N-1]\cup[N+1,\infty)$ into $\mathbb{C}$ connecting the points $(m,\Theta(m))$ where $m\in\mathbb{Z}\cap\left((-\infty,-N-1]\cup[N+1,\infty)\right).$ More precisely, on each segment [m1,m][m-1,m] with $m\in\mathbb{Z}\cap((-\infty,-N-1]\cup[N+2,\infty)),$ the function $\tilde{\Theta}$ is given as
@@
[…]
@@
Consider the weighted composition operator $\tilde{T}_{\alpha,w}$ on $C_{0}(\mathbb{R})$ given by $\tilde{T}_{\alpha,w}(f)=w\cdot(f\circ\alpha),$ where $0<w,w^{-1}\in C_{b}(\mathbb{R}),$ and $\alpha$ is a homeomorphism of $\mathbb{R}.\\$ If $\lim_{n\to\infty}\prod_{k=1}^{n}(w\circ\alpha^{-k})^{-1}(n)=0,$ then the set
$\left\{f\in C_{0}(\mathbb{R}):\|\tilde{T}_{\alpha,w}^{n}(f)\|_{\infty}\geq 1\text{ for all }n\in\mathbb{N}\right\}\\$
[…]
@@
$(t-n)\left(\prod_{k=1}^{n+1}(w\circ\alpha^{-k})^{-1}(n+1)-\prod_{k=1}^{n}(w\circ\alpha^{-k})^{-1}(n)\right)$
for $t\in[n,n+1],$ whereas for $t\in[0,1],$ we put $g(t)=t(w\circ\alpha^{-k})^{-1}(1).$ Moreover, we let $g(t)=0$ for $t\leq 0.$ It is easily seen that $g\geq 0$ and $g\in C_{0}(\mathbb{R})$ since $\lim_{n\to\infty}\prod_{k=1}^{n}(w\circ\alpha^{-k})^{-1}(n)=0$ by the assumption. By Theorem 5.3 , the set $\Gamma_{g}$ is not $\sigma\text{-porous.}$ Now, for each $f\in\Gamma_{g}$ and $n\in\mathbb{N},$ we have that